Place in standard form for the circle ( which is (x-h)^2 + (y-k)^2 = r^2):
x^2 + y^2 + 8x - 14y + 59 = 0

Answer by ewatrrr(24785)

 
Hi,
x^2 + y^2 + 8x - 14y + 59 = 0
completing the respective Squares...
(x^2+8x +16) + (y^2 -14y + 49) -16-49 + 59 = 0
(x+4)^2 + (y-7)^2 = 6